# Unmasking real-world audio deepfakes: A data-centric approach

### v1

@@ 1 Introduction
“As you know our FTX exchange is going bankrupt, but you should not panic[…] we have prepared a giveaway. Just go to the site …” We are assured by Sam Bankman-Fried (FTX CEO) in a viral clip posted on Twitter in November 2022. Or are we? The video turned out to be fake, but many were fooled by it and, consequently, lost money. Such automatically generated clips (known as deepfakes) are permeating our society resulting in misinformation, scams and making us doubt the veracity of what we are seeing daily on the internet.
@@ 4.1 Scientific dataset selection
In Section 3 we showed that when it comes to more recent deepfake samples (i.e., the AI4T dataset), the contents of ASV19 is unable to provide a good generalisation performance for the underlying model. However, as seen in Section 2.1, there are many other datasets that could be used for training. In this section we aim to find the combination that yields the best results. We consider the seven scientific datasets and evaluate all $128$ ($2^{7}$) combinations. For each dataset we use all of its samples, including real samples and those in the test splits. We use layer 9 features and no data augmentation.
@@ 4.2 Sample pruning
Although the deepfake datasets were released as cohesive units, not all their samples may be equally relevant for the generalisation objective. As a result, similar to [31, 17, 32], we explore three data pruning strategies: (i) random pruning; (ii) data-informed pruning; and (iii) algorithm-informed pruning. Random pruning refers to randomly discarding samples from the training data, which was shown to exhibit good generalisation when no sample scoring can used [33, 34]. For data-informed pruning, we rank samples by the distance to their mean, and select either the closest (cluster-closest) or the furthest (cluster-furthest) samples. The samples are represented by their average pooled self-supervised embeddings and compared using the Euclidean distance. We apply this strategy independently on both real and fake samples from each of $N$ datasets, and obtain $2N$ sets of samples, which we ensemble in the final pruned dataset. For algorithm-informed pruning, we use the logistic regression’s margins over the samples. We remove the closest (margin-noisy) or the closest and furthest (margin-both) points with respect to the decision hyperplane, irrespective of the dataset that they belong to.
Figure 4 shows a subset of the results of the sample pruning methods (complete results are available in the code repository). We use either the best combination of datasets (i.e. For+ODSS+MLAAD) or the complete set of available samples (ALL data) as the starting pool for the pruning strategies, and report the EER over the real-world datasets. First, we observe that performance is relatively stable with the amount of discarded data. This means that many of the samples carry redundant information. For the data pruned from the FoR+ODSS+MLAAD combination (left side of the figure), the average results obtained after pruning are similar to the baseline. The only pruning method that performs marginally better is margin-both. However, obtaining a similar performance when discarding 90% of the data is in itself a very important observation, enabling a more efficient model selection. When pruning the complete set of samples (right side of the figure), margin-based selections work best for both ITW and AI4T. For AI4T, almost all strategies halve the baseline EER (bottom right plot). This may be because the baseline for AI4T is rather poor, at 29.03% EER, and any noisy and outlier samples’ reduction are relevant to the logistic regressor.

### v2

@@ 1 Introduction
``As you know our FTX exchange is going bankrupt, but you should not panic[…] we have prepared a giveaway. Just go to the site …'' We are assured by Sam Bankman-Fried (FTX CEO) in a viral clip posted on Twitter in November 2022. Or are we? The video turned out to be fake, but many were fooled by it and, consequently, lost money. Such automatically generated clips (known as deepfakes) are permeating our society resulting in misinformation, scams and making us doubt the veracity of what we are seeing daily on the internet.
@@ 4.1 Scientific dataset selection
In Section 3 we showed that when it comes to more recent deepfake samples (i.e., the AI4T dataset), the contents of ASV19 is unable to provide a good generalisation performance for the underlying model. However, as seen in Section 2.1, there are many other datasets that could be used for training. In this section we aim to find the combination that yields the best results. We consider the seven scientific datasets and evaluate all 128128 ($2^{7}$) combinations. For each dataset we use all of its samples, including real samples and those in the test splits. We use layer 9 features and no data augmentation.
@@ 4.2 Sample pruning
Although the deepfake datasets were released as cohesive units, not all their samples may be equally relevant for the generalisation objective. As a result, similar to [31, 17, 32], we explore three data pruning strategies: (i) random pruning; (ii) data-informed pruning; and (iii) algorithm-informed pruning. Random pruning refers to randomly discarding samples from the training data, which was shown to exhibit good generalisation when no sample scoring can used [33, 34]. For data-informed pruning, we rank samples by the distance to their mean, and select either the closest (cluster-closest) or the furthest (cluster-furthest) samples. The samples are represented by their average pooled self-supervised embeddings and compared using the Euclidean distance. We apply this strategy independently on both real and fake samples from each of $N$ datasets, and obtain $2N$ sets of samples, which we ensemble in the final pruned dataset. For algorithm-informed pruning, we use the logistic regression's margins over the samples. We remove the closest (margin-noisy) or the closest and furthest (margin-both) points with respect to the decision hyperplane, irrespective of the dataset that they belong to.
Figure 4 shows a subset of the results of the sample pruning methods (complete results are available in the code repository). We use either the best combination of datasets (i.e. For+ODSS+MLAAD) or the complete set of available samples (ALL data) as the starting pool for the pruning strategies, and report the EER over the real-world datasets. First, we observe that performance is relatively stable with the amount of discarded data. This means that many of the samples carry redundant information. For the data pruned from the FoR+ODSS+MLAAD combination (left side of the figure), the average results obtained after pruning are similar to the baseline. The only pruning method that performs marginally better is margin-both. However, obtaining a similar performance when discarding 90% of the data is in itself a very important observation, enabling a more efficient model selection. When pruning the complete set of samples (right side of the figure), margin-based selections work best for both ITW and AI4T. For AI4T, almost all strategies halve the baseline EER (bottom right plot). This may be because the baseline for AI4T is rather poor, at 29.03% EER, and any noisy and outlier samples' reduction are relevant to the logistic regressor.